OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5 over 9 (F − 32), where F stands for degrees Fahrenheit. Part 1: Solve the equation for F and show all steps. Part 2: Determine how many degrees Fahrenheit 15 degrees Celsius is.

OpenStudy (anonymous):

C=9/5(f-32) 5/9C=F-32 F=5/9C+32 6=5/9c+32 -26=5/9c c=-46.8 degrees celsius

OpenStudy (anonymous):

Convert 15 degrees celsius to farenheit.

OpenStudy (anonymous):

=15*9/5+32 =59 degrees farenheit

OpenStudy (anonymous):

C=5/9(F-32) This is the correct equation. F=C*9/5+32 this is the correct equation to get F. 15 C is 59 F